OpenStudy (ashley1nonly):

Rearranging equation>>>>>The first-order decomposition of cyclopropane has a rate constant of 6.7 x 10-4 s-1. If the initial concentration of cyclopropane is 1.33 M, what is the concentration of cyclopropane after 644 s?

OpenStudy (ashley1nonly):

i know that im using the first order reaction which is ln ([A]t/[A]o=-kt we are solving for [A]t but how would i get rid of the ln so that i can solve for it???

OpenStudy (aaronq):

\(\large ln(\dfrac{[A]_t}{[A]_0})=-kt\) \(\large e^{ln(\frac{[A]_t}{[A]_0})}=e^{-kt}\) \(\large \dfrac{[A]_t}{[A]_0}=e^{-kt}\)